Two plastic balls, 1 and 2 are rubbed together with a cloth until they both carry a uniformly disturbed charge of magnitude 0.10 NC. Ball 1 is negatively charged and ball 2 is positively charged. The balls are held apart by 900mm stick that turns to the centres of both balls


What is the magnitude of the dipole moment



1
Expert's answer
2021-09-03T08:25:07-0400

The electric dipole moment is given by

"p=qd\\=(0.1\times 10^{-9}\:\rm C)\times(0.9\: m)=9\times 10^{-11}\: C\cdot m"